OpenStudy (anonymous):

Two trains leave stations 182 miles apart at the same time and travel toward each other. One train travels at 55 miles per hour while the other travels at 75 miles per hour. How long will it take for the two trains to meet?

OpenStudy (anonymous):

(distance apart)/(speed1+speed2) see if it works

OpenStudy (anonymous):

1.4 hours

OpenStudy (anonymous):

awesome you were right with that formula!! thank ya

OpenStudy (anonymous):

nice.. i just learned figured that out today.. id like to make sure it works all the time

OpenStudy (anonymous):

do you know the formula to find rate?

OpenStudy (anonymous):

rate of what? distance/time ?

OpenStudy (anonymous):

Two cars leave towns 510 kilometers apart at the same time and travel toward each other. One car's rate is 12 kilometers per hour less than the other's. If they meet in 3 hours, what is the rate of the slower car?

OpenStudy (anonymous):

oh same equation.. you just need to rearrange it and use a variable for one of the speeds

OpenStudy (anonymous):

510/12+b

OpenStudy (anonymous):

if you do it that way.. then think of it as (12+(b-12))

OpenStudy (anonymous):

so 12(b-12)=510?

OpenStudy (anonymous):

im way off on this one

OpenStudy (anonymous):

ill rearrange it for you so you can learn too.. d/(s1+s2) = t so d/t = s1 + s2

OpenStudy (anonymous):

... s2 = s1-12

OpenStudy (anonymous):

substituting ... d/t = s1 + (s1-12) d/t = 2s1 -12

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

yes it did thank you doing more on rates soo seeing if i can keep getting them right

OpenStudy (anonymous):

510/3 = s + s-12 170 = 2s - 12 182 = 2s 182/2 = s 91 = s s-12 = slower car 91 - 12 = slower car

OpenStudy (anonymous):

is that the correct answer.. im practicing too.. i by no means am any expert.. i just want to get better and this makes sense to me

OpenStudy (anonymous):

yeah that formula you gave me was correct

OpenStudy (anonymous):

Ravi hiked up a mountain trail and camped overnight at the top. The next day he returned down the same trail. His average rate traveling uphill was 2.4 kilometers per hour and his average rate downhill was 3.6 kilometers per hour. If he spent a total of 11 hours hiking, how long was the trail?

OpenStudy (anonymous):

same thing for this one?

OpenStudy (anonymous):

like 3.6+2.4= 11/?

OpenStudy (anonymous):

rearrange the formula and see if you get the right answer bc thats what i would have done

OpenStudy (anonymous):

oops d= (s1+s2)t

OpenStudy (anonymous):

did you get 66km?

OpenStudy (anonymous):

or is that answer right?

OpenStudy (anonymous):

no thats not the right answer

OpenStudy (anonymous):

hmm. what is the correct answer?

OpenStudy (anonymous):

it was like 15 something i wasnt able to get that one either

OpenStudy (anonymous):

d=rt where r = rate and d = distance and t = time here is your equation.. d = t(r1*r2)/(r1+r2)